OpenStudy (anonymous):

In 2003, online holiday sales were $17 billion, and in 2006, they were $26 billion. (a) Find a linear function S that models these data, where x is the year. (b) Interpret the slope of the graph of S. (c) Predict when online holiday sales might reach $41 billion.

OpenStudy (anonymous):

(a) point A (2003, 17) point B (2006, 26) \[slope = { {\Delta y} \over { \Delta x } } = { { 26 - 17 } \over { 2006 - 2003 } } = { 9 \over 3 } = 3\] \[s(x)= ax + b\] Now substitute slope and 1 point to find b \[26 = 3 \times 2006 + b \implies b = 26 - 6018 = -5992\] \[s(x) = 3x -5992\] (b) The slope is the amount (in billions) with which the holiday sales increase each year (c) \[41 = 3x - 5992 \implies x = 2011\]

OpenStudy (anonymous):

Thank you so much sir, very well done!